OpenStudy (anonymous):

how to foil 4x^2-9X+2

OpenStudy (anonymous):

factor using FOIL (4x-1)(x-2)=0 4x-1=0 4x=1 x=1/4 or x-2=0 x=2

OpenStudy (jack1):

sorry, do over, missed the negative before \(\Large 4x^2-9x+2 \) \(\Large (4x - 1)(x-2) \)

OpenStudy (anonymous):

that is all i new sorry :(

OpenStudy (jack1):

think that's right... First: 4x * x = 4x^2 Outer: 4x * -2 = -8x Inner: -1 * x = -1x Last: -1 * -2 = 2 add like terms 4x^2 -8x - 1x +2 = 4x^2 -9x +2

OpenStudy (jack1):

so if it =0, what you said above is perfectly right 4x - 1 = 0, so x = 1/4 or x-2 = 0, so x = 2

terenzreignz (terenzreignz):

Nobody likes to be foiled :>

OpenStudy (jack1):

;o lol, heya TJ ! ;)

terenzreignz (terenzreignz):

Hello :)